# Adaptive Gradient Normalization and Independent Sampling for (Stochastic) Generalized-Smooth Optimization

### v2

@@ 4 Generalized-Smooth nonconvex Optimization
In the following sections, we first consider deterministic generalized-smooth optimization and study the impact of adaptive gradient normalization on the convergence rate of gradient methods. Then, we consider stochastic generalized-smooth optimization and propose a novel independent sampling scheme for improving the convergence guarantee of stochastic gradient methods.
@@ 5 Adaptive Gradient Normalization for Deterministic Generalized-Smooth Optimization
In deterministic generalized-smooth optimization, many previous works have empirically demonstrated the faster convergence of normalized gradient descent-type algorithms (e.g., clipped GD) over the standard gradient descent algorithm in various machine learning applications (Chen et al., 2023; Zhang et al., 2019). On the other hand, theoretically, these algorithms were only shown to achieve the same iteration complexity $\mathcal{O}(\epsilon^{-2})$ as the gradient descent algorithm in generalized-smooth optimization. In this section, to further advance the theoretical understanding and explain the inconsistency between theory and practice, we explore the advantage of adapting gradient normalization to the special Polyak-Łojasiewicz-type (PŁ) geometry in generalized-smooth optimization. We aim to show that gradient normalization, when properly adapted to the underlying PŁ geometry, can help accelerate the convergence rate in generalized-smooth optimization.
@@
The above generalized PŁ condition is inspired by the Kuradyka-Łojasiewicz (KŁ)-exponent condition proposed in (Li & Pong, 2018). When $\rho>1$, equation 5 reduces to the KŁ-exponent condition. When $\rho=2$, equation 5 reduces to the standard PŁ condition. Moreover, some recent works have shown that PŁ-type geometries widely exist in the loss landscape of phase retrieval (Zhou et al., 2016) and over-parametrized deep neural networks (Liu et al., 2022a; Scaman et al., 2022) etc, and we hope that our analysis based on Assumption 2 will allow researchers to rethink the relationship between adaptive normalization and loss landscape geometry.
@@
Theorem 1 (Convergence Rate of AN-GD)
Let Assumptions 1 and 2 hold. Denote $\Delta_{t}:=f(\omega_{t})-f^{*}$ as the function value gap. Choose target error $0<\epsilon\leq\min\{1,\frac{1}{2\mu}\}$ and define learning rate $\gamma=\frac{(2\mu\epsilon)^{\beta/\rho}}{8(L_{0}+L_{1})+1}\leq\frac{2}{\mu}$ for some $\beta\in[\alpha,1]$. Then, the following statements hold,
  • If $\beta<2-\rho$, in order to achieve $\Delta_{T}\leq\epsilon$, the total number of iteration must satisfy
    $T\geq\max\Big{\{}\frac{8\rho(8(L_{0}+L_{1})+1)}{(2-\beta-\rho)(2\mu)^{2/\rho}\epsilon^{(2-\rho)/{\rho}}},\frac{1}{(2^{(2-\beta-\rho)/(2-\beta)}-1)(\Delta_{0}\epsilon)^{(2-\beta-\rho)/{\rho}}}\Big{\}}=\Omega\big{(}(\frac{1}{\epsilon})^{\frac{2-\rho}{\rho}}\big{)}.$ (7)
  • If $\beta=2-\rho$, in order to achieve $\Delta_{T}\leq\epsilon$, the total number of iteration must satisfy
@@
  • If $1\geq\beta>2-\rho$, and for some $T_{0}$ satisfying $\Delta_{T_{0}}\leq\mathcal{O}((\gamma\mu)^{\frac{\rho}{\rho+\beta-2}})$, then the total number of iterations after $T_{0}$ must satisfy
    $T\gtrsim$ $\Omega\Big{(}\log\Big{(}\big{(}\frac{1}{\epsilon}\big{)}^{\frac{\beta}{\rho+\beta-2}}\Big{)}+\log\Big{(}\log\Big{(}\frac{(2\mu)^{(2+\beta)/\rho}}{4(8(L_{0}+L_{1})+1)^{2}\epsilon}\Big{)}\Big{)}\Big{)}=\Omega\big{(}\log\big{(}(\frac{1}{\epsilon})^{\frac{\beta}{\rho+\beta-2}}\big{)}\big{)}.$ (9)
Theorem 1 characterizes the convergence rates of AN-GD under different choices of $\rho$ and $\beta$. In particular, when $\alpha=\beta=0$ and $\rho=2$, Theorem 1 reduces to the linear convergence rate achieved by gradient descent under the standard PŁ and $L$-smooth condition. Theorem 1 also guides the choice of gradient normalization parameter $\beta$ under different geometric conditions. For example, if there exists $\beta\in[\alpha,1]$ such that $\rho+\beta>2$, AN-GD can boost its convergence rate from polynomial to local linear convergence. It also indicates when $\rho$ is very small such that $\rho+\beta<2$, the effects of $\beta$ can be marginal. However, for $\rho+\beta\leq 2$, the iteration complexities depend more on the specific values of $\rho$ and $\beta$. For example, when $\rho=1$ and consider two different choices of gradient normalization parameters $\beta_{1}=\frac{2}{3},\beta_{2}=1$, AN-GD achieves the iteration complexities $\mathcal{O}((\frac{1}{\epsilon}))$ and $\mathcal{O}(\frac{1}{\epsilon}\log(\frac{1}{\epsilon}))$, respectively. This result matches the empirical observation made in (Chen et al., 2023) that choosing a smaller $\beta\in[\alpha,1]$ can improve the convergence rate.
We note that several concurrent works also studied normalized gradient descent methods in deterministic convex settings. In Gorbunov et al. (2024), the authors consider various update rules such as $w_{t+1}=w_{t}-\gamma\tfrac{\nabla f(w_{t})}{L_{0}+L_{1}\|\nabla f(w_{t})\|},\quad\text{and}\quad w_{t+1}=w_{t}-\gamma\tfrac{f(w_{t})-f(w^{*})}{\|\nabla f(w_{t})\|}\nabla f(w_{t}).$ For generalized smooth convex optimization, they obtain a two-phase convergence result, similar to the case when $1\geq\beta>2-\rho$ in Theorem 1, where the second phase achieves complexity bound $\mathcal{O}(\epsilon^{-1})$. They further employ the triangle method to derive a global $\mathcal{O}(\epsilon^{-1})$ complexity bound. In Vankov et al. (2024b), the authors propose a related framework by choosing $\gamma=\frac{1}{\sqrt{T}\|\nabla f(w_{t})\|}$, which aligns with our AN-GD when $\beta=1$. They also obtain $\mathcal{O}(\epsilon^{-1})$ complexity for generalized smoothness convex optimization.
@@
In Li et al. (2024), it has been shown that GD achieves $\mathcal{O}(\log(\frac{1}{\epsilon}))$, $\mathcal{O}({\epsilon}^{-1})$, $\mathcal{O}({\epsilon^{-2}})$ complexity bounds for generalized-smooth strongly convex, convex and non-convex optimization, respectively. Compared with Chen et al. (2023), both AN-GD and GD achieve $\mathcal{O}({\epsilon^{-2}})$ complexity bound for generalized non-convex optimization, and hence are optimal deterministic methods. On the other hand, the following Proposition 1 further obtains the convergence rate of gradient descent (GD) under the same Assumptions 1 and 2. See Proof 5 in Appendix F.
@@
Let Assumptions 1 and 2 hold. Assume there exists a positive constant $G$ such that $\|\nabla f(x_{t})\|\leq G,\forall t\in T$, When $\rho=\alpha=1$ and setting $\gamma\leq\min\{\frac{1}{L_{0}},\frac{1}{2L_{1}G}\}$, gradient descent converges to an $\epsilon$-stationary point within $T=\Omega(\frac{G}{\epsilon})$ iterations.
From Theorem 1, under the setting $\rho=\alpha=1$ and with $\beta=1$, AN-GD converges to an $\epsilon$-stationary point in $\mathcal{O}(\frac{1}{\epsilon}\log(\frac{1}{\epsilon}))$ iterations. However, the convergence guarantee for GD requires an additional assumption that the gradient norm is upper bounded by a constant $G$. This constant typically depends on the function value gap $\Delta_{0}$ and $\|x_{0}-x^{*}\|$, which can be large in generalized smooth optimization settings. Such large value can be the key reason for the slower empirical convergence of GD observed in (Chen et al., 2023).
@@ 6.1 Normalized SGD and Its Limitations
AN-SGD-type algorithms have attracted a lot of attention recently for solving stochastic generalized-smooth problems (Zhang et al., 2019; 2020a; Liu et al., 2022b; Jin et al., 2021). In particular, previous works have shown that, when choosing $\beta=1$ and using gradient clipping or momentum acceleration, AN-SGD’s variations (e.g., NSGD with momentum (Cutkosky & Mehta, 2020; Jin et al., 2021), clipped SGD (Zhang et al., 2019; 2020a)) can achieve a sample complexity of $\mathcal{O}(\epsilon^{-4})$. This result matches that of the standard SGD for solving classic stochastic smooth problems (Ghadimi & Lan, 2013). However, AN-SGD has several limitations as summarized below.
@@
  • Strong assumption: To control the estimation bias and establish theoretical convergence guarantee for ANSGD-type algorithms in generalized-smooth nonconvex optimization, the existing studies need to adopt strong assumptions. For example, Zhang et al. (2019; 2020a) and Liu et al. (2022b) assume that the stochastic approximation error $\|\nabla f_{\xi}(w)-\nabla F(w)\|$ is bounded by a constant almost surely. In practical applications, this constant can be a large numerical number if certain sample $\xi$ happen to be an outlier. On the other hand, Chen et al. (2023); Reisizadeh et al. (2023) require ANSGD-type of algorithms such as SPIDER (Fang et al., 2018) to have $\mathcal{O}(\epsilon^{-2})$ batch size to ensure fast convergence. It essentially makes it a full-batch algorithm and is impractical in practice.
@@ 6.2 Independently-and-Adaptively Normalized SGD
where $\xi$ and $\xi^{\prime}$ are samples draw independently from the underlying data distribution. Intuitively, the independence between $\xi$ and $\xi^{\prime}$ decorrelates the denominator from the numerator, making it an unbiased stochastic gradient estimator (up to a scaling factor). Specifically, we formally have that
@@ 6.3 Convergence Analysis of IAN-SGD
We adopt the following standard assumptions on the stochastic gradient.
@@
Assumption 4 (Approximation error )
@@
We note that the above Assumption 4 is weaker than the bounded approximation error assumption (i.e., $\tau_{1}=0$) adopted in (Zhang et al., 2019; 2020a; Liu et al., 2022b). Specifically, it allows the approximation error to scale with the full gradient norm and only assumes bounded error at the stationary points. With these assumptions, we establish the following upper bounds regarding $\|\nabla F(w_{t})\|$, $\|\nabla f_{\xi}(w_{t})\|$. See Proof 6 in Appendix G.
@@
In the existing work (Zhang et al., 2019), they studied the iteration complexity of clipped SGD (i.e., $\beta=1$ and $\xi^{\prime}=\xi$, without independence sampling) in the special bounded noise setting $\tau_{1}=0$, and obtained the complexity
$\Lambda\max\Big{\{}\frac{4\Lambda}{\epsilon^{4}},\frac{128L_{1}}{\epsilon},\frac{80L_{0}+512L_{1}\tau_{2}}{\epsilon^{2}}\Big{\}},$ (18)
@@
In the existing work Zhang et al. (2020a), they analyze a generalized clipping framework that combines momentum acceleration and gradient clipping, namely $w_{t+1}=w_{t}-\big{(}\nu\min\big{\{}c,\tfrac{\gamma}{\|m_{t}\|}\big{\}}m_{t}+(1-\nu)\min\big{\{}c,\tfrac{\gamma}{\|\nabla f_{\xi}(w_{t})\|}\big{\}}\nabla f_{\xi}(w_{t})\big{)},$ under a bounded noise setting (i.e., $\tau_{1}=0$). They establish a complexity of $\mathcal{O}\big{(}(F(w_{0})-F^{*})\tau_{2}^{2}L_{0}/\epsilon^{4}\big{)}$ when $\gamma/c=\Theta(\tau_{2})$, where $m_{t}$ denotes the momentum term and $c$ is the clipping threshold. This framework enables them to construct a recursion based directly on $\|\nabla F(w_{t})\|$, whereas our analysis is built around recursive bounds involving $\|\nabla F(w_{t})\|^{2}$ or $\|\nabla F(w_{t})\|^{2-\beta}$ at each step. Consequently, while both approaches yield the same order of convergence in terms of $\epsilon$, the complexity expressions slightly differ in the numerators due to different recursions.
Compared to the existing studies (Zhang et al., 2019; 2020b; Reisizadeh et al., 2023; Zhang et al., 2020a; Koloskova et al., 2023b) on gradient clipping algorithms, Theorem 2 establishes convergence guarantee without the bounded approximation error assumption and the use of extremely large batch sizes that depend on the target error $\epsilon$. Through numerical experiments in Section 7 and ablation study in A.2 later, we show that it suffices to query a small number of independent samples for IAN-SGD in practice.

### v3

@@ 4 Generalized-Smooth nonconvex Optimization
In the following sections, we first consider deterministic generalized-smooth optimization and study the impact of adaptive gradient normalization on the convergence rate of gradient methods. Then, we consider stochastic generalized-smooth optimization and propose a novel independent sampling scheme for improving the convergence guarantee of stochastic gradient methods under relaxed noise assumption.
@@ 5 Adaptive Gradient Normalization for Deterministic Generalized-Smooth Optimization
In deterministic generalized-smooth optimization, several previous works have empirically demonstrated the faster convergence of normalized gradient descent-type algorithms (e.g., clipped GD) over the standard gradient descent algorithm in various machine learning applications (Zhang et al., 2019; Chen et al., 2023). On the other hand, theoretically, these algorithms were only shown to achieve the same iteration complexity $\mathcal{O}(\epsilon^{-2})$ as the gradient descent algorithm in generalized-smooth optimization. In this section, to further advance the theoretical understanding and explain the inconsistency between theory and practice, we explore the advantage of adapting gradient normalization to the special Polyak-Łojasiewicz-type (PŁ) geometry in generalized-smooth optimization. We aim to show that gradient normalization, when properly adapted to the underlying PŁ geometry, can help accelerate the convergence rate in generalized-smooth optimization.
@@
The above generalized PŁ condition is inspired by the Kurdyka-Łojasiewicz (KŁ)-exponent condition proposed in Li & Pong (2018). When $\rho>1$, equation 5 reduces to the KŁ-exponent condition. When $\rho=2$, equation 5 reduces to the standard PŁ condition. Moreover, some recent works have shown that PŁ-type geometries widely exist in the loss landscape of phase retrieval (Zhou et al., 2016) and over-parametrized deep neural networks (Liu et al., 2022a; Scaman et al., 2022), and we hope that our analysis based on Assumption 2 will allow researchers to rethink the relationship between adaptive normalization and loss landscape geometry.
@@
Theorem 1 (Convergence of AN-GD)
Let Assumptions 1 and 2 hold. Denote $\Delta_{t}:=f(w_{t})-f^{*}$ as the function value gap. Choose target error $0<\epsilon\leq\min\{1,\frac{1}{2\mu}\}$ and define learning rate $\gamma=\frac{(2\mu\epsilon)^{\beta/\rho}}{8(L_{0}+L_{1})+1}\leq\frac{2}{\mu}$ for some $\beta\in[\alpha,1]$. Then, the following statements hold,
  • If $\beta<2-\rho$, in order to achieve $\Delta_{T}\leq\epsilon$, the total number of iterations must satisfy
    $T\geq\max\Big{\{}\frac{8\rho(8(L_{0}+L_{1})+1)}{(2-\beta-\rho)(2\mu)^{2/\rho}\epsilon^{(2-\rho)/{\rho}}},\frac{1}{(2^{(2-\beta-\rho)/(2-\beta)}-1)\Delta_{0}^{{(\rho+\beta-2)}/{\rho}}\epsilon^{(2-\beta-\rho)/{\rho}}}\Big{\}}=\Omega\big{(}(\frac{1}{\epsilon})^{\frac{2-\rho}{\rho}}\big{)}.$ (7)
  • If $\beta=2-\rho$, in order to achieve $\Delta_{T}\leq\epsilon$, the total number of iterations must satisfy
@@
  • If $1\geq\beta>2-\rho$, and for some $T_{0}$ satisfying $\Delta_{T_{0}}\leq\mathcal{O}\big{(}(\gamma\mu^{(2-\beta)/(\rho+\beta-2)})^{\frac{\rho}{\rho+\beta-2}}\big{)}$, then the total number of iterations after $T_{0}$ must satisfy
    $T\gtrsim$ $\Omega\Big{(}\log\Big{(}\big{(}\frac{1}{\epsilon}\big{)}^{\frac{\beta}{\rho+\beta-2}}\Big{)}+\log\Big{(}\log\Big{(}\frac{(2\mu)^{2/(\rho+\beta-2)}}{(32(L_{0}+L_{1})+4)^{\rho/(\rho+\beta-2)}\epsilon}\Big{)}\Big{)}\Big{)}=\Omega\big{(}\log\big{(}(\frac{1}{\epsilon})^{\frac{\beta}{\rho+\beta-2}}\big{)}\big{)}.$ (9)
Theorem 1 characterizes the convergence rates of AN-GD under different choices of $\rho$ and $\beta$. In particular, when $\alpha=\beta=0$ and $\rho=2$, Theorem 1 reduces to the linear convergence rate achieved by gradient descent under the standard PŁ and $L$-smooth condition. Theorem 1 also guides the choice of gradient normalization parameter $\beta$ under different geometric conditions. For example, if there exists $\beta\in[\alpha,1]$ such that $\rho+\beta>2$, AN-GD can boost its convergence rate from polynomial to local linear convergence. It also indicates when $\rho$ is very small such that $\rho+\beta<2$, the effects of $\beta$ can be marginal. However, for $\rho+\beta\leq 2$, the iteration complexities depend more on the specific values of $\rho$ and $\beta$. For example, when $\rho=1$ and consider two different choices of gradient normalization parameters $\beta_{1}=\frac{2}{3},\beta_{2}=1$, AN-GD achieves the iteration complexities $\mathcal{O}({\epsilon}^{-1})$ and $\tilde{\mathcal{O}}({\epsilon}^{-1})$, respectively. This result matches the empirical observation made in (Chen et al., 2023) that choosing a smaller $\beta\in[\alpha,1]$ can improve the convergence rate.
We note that several concurrent works also studied normalized gradient descent methods in deterministic convex settings. In Gorbunov et al. (2024), the authors consider various update rules such as $w_{t+1}=w_{t}-\gamma\tfrac{\nabla f(w_{t})}{L_{0}+L_{1}\|\nabla f(w_{t})\|}\quad\text{and}\quad w_{t+1}=w_{t}-\gamma\tfrac{f(w_{t})-f(w^{*})}{\|\nabla f(w_{t})\|}\nabla f(w_{t}).$ For generalized smooth convex optimization, they obtain a two-phase convergence result, similar to the case when $1\geq\beta>2-\rho$ in Theorem 1, where the second phase achieves complexity bound $\mathcal{O}(\epsilon^{-1})$. They further employ the triangle method to derive a global $\mathcal{O}(\epsilon^{-1})$ complexity bound. In Vankov et al. (2024b), the authors propose a related framework by choosing $\gamma=\frac{1}{\sqrt{T}\|\nabla f(w_{t})\|}$, which aligns with our AN-GD when $\beta=1$. They also obtain $\mathcal{O}(\epsilon^{-1})$ complexity for generalized smoothness convex optimization.
@@
In Li et al. (2024), it has been shown that GD achieves $\mathcal{O}(\log({\epsilon}^{-1}))$, $\mathcal{O}({\epsilon}^{-1})$, $\mathcal{O}({\epsilon^{-2}})$ complexity bounds for generalized-smooth strongly convex, convex and non-convex optimization, respectively. Compared with Chen et al. (2023), both AN-GD and GD achieve $\mathcal{O}({\epsilon^{-2}})$ complexity bound for generalized non-convex optimization, and hence are optimal deterministic methods. On the other hand, the following Proposition 1 further obtains the convergence rate of gradient descent (GD) under the same Assumptions 1 and 2. See Proof 5 in Appendix F.
@@
Let Assumptions 1 and 2 hold. Assume there exists a positive constant $G$ such that $\|\nabla f(x_{t})\|\leq G,\forall t\in T$, When $\rho=\alpha=1$ and setting $\gamma\leq\min\{\frac{1}{L_{0}},\frac{1}{2L_{1}G}\}$, gradient descent converges to an $\epsilon$-stationary point within $T\geq\Omega(\frac{G}{\epsilon})$ iterations.
From Theorem 1, under the setting $\rho=\alpha=1$ and with $\beta=1$, AN-GD converges to a $\epsilon$-stationary point in $\tilde{\mathcal{O}}({\epsilon}^{-1})$ iterations. However, the convergence guarantee for GD requires an additional assumption that the gradient norm is upper bounded by a constant $G$. This constant typically depends on the function value gap $\Delta_{0}$ and $\|x_{0}-x^{*}\|$, which can be large in generalized smooth optimization settings. Such large value can be the key reason for the slower empirical convergence of GD observed in Chen et al. (2023).
@@ 6.1 Normalized SGD and Its Limitations
AN-SGD-type algorithms have attracted a lot of attention recently for solving stochastic generalized-smooth problems (Zhang et al., 2019; 2020a; Liu et al., 2022b; Jin et al., 2021). In particular, previous works have shown that, when choosing $\beta=1$ and using gradient clipping or momentum acceleration, AN-SGD’s variations (e.g., NSGD with momentum (Cutkosky & Mehta, 2020; Jin et al., 2021), clipped SGD (Zhang et al., 2019; 2020a)) can achieve a sample complexity of $\mathcal{O}(\epsilon^{-4})$. This result matches the convergence order of the standard SGD (Ghadimi & Lan, 2013) for solving classic stochastic smooth problems. However, AN-SGD has several limitations as summarized below.
@@
  • Strong assumptions: To control the estimation bias and establish theoretical convergence guarantee for ANSGD-type algorithms in generalized-smooth nonconvex optimization, the existing studies need to adopt strong assumptions. For example, Zhang et al. (2019; 2020a) and Liu et al. (2022b) assume that the stochastic approximation error $\|\nabla f_{\xi}(w)-\nabla F(w)\|$ is bounded by a constant almost surely. In practical applications, this constant can be a large numerical number if certain sample $\xi$ happen to be an outlier. On the other hand, Chen et al. (2023); Reisizadeh et al. (2023) require AN-SGD type of algorithms such as SPIDER (Fang et al., 2018) to have $\Omega(\epsilon^{-2})$ batch size to ensure fast convergence. It essentially makes it a full-batch algorithm and is impractical in practice.
@@ 6.2 Independently-and-Adaptively Normalized SGD
where $\xi$ and $\xi^{\prime}$ are samples draw independently from the underlying data distribution. Intuitively, the independence between $\xi$ and $\xi^{\prime}$ decorrelates the denominator from the numerator, making it an unbiased stochastic gradient estimator (up to a scaling factor). Specifically, we have that
@@ 6.3 Convergence Analysis of IAN-SGD
We adopt the following bias assumption on the stochastic gradient $\nabla f_{\xi}(w)$.
@@
Assumption 4 (Affine approximation Error)
@@
Assumption 4 allows the approximation error of stochastic gradient scaling with the full gradient norm on a per-sample basis, while only requiring the error to remain bounded at stationary points. Compared to other noise assumptions, Assumption 4 is much weaker than the almost sure error bound (i.e., $\tau_{1}=0$) adopted in Zhang et al. (2019; 2020a); Liu et al. (2022b) but stronger than expected error bound (i.e., bounded variance assumption $\mathbb{E}_{\xi}\|\nabla f(w)-\nabla F(w)\|^{2}\leq\tau_{2}^{2}$) adopted in Koloskova et al. (2023); Reisizadeh et al. (2023). Our motivation for studying the convergence of IAN-SGD under this noise assumption is to strike a balance between achieving standard convergence rates (i.e., $\mathcal{O}(\epsilon^{-4})$) under relaxed noise conditions, while keeping the batch size at the $\Omega(1)$-level. As later we show in Theorem 2, IAN-SGD achieves $\mathcal{O}(\epsilon^{-4})$ sample complexity under Assumption 4.
Remark 3
Existing work (Koloskova et al., 2023) has shown that when adopting $\Omega(1)$ sample per iteration, clipped SGD attains $\mathcal{O}(\epsilon^{-5})$ sample complexities under generalized-smooth and bounded variance assumptions. In order to achieve the standard $\mathcal{O}(\epsilon^{-4})$ sample complexity under such assumption, Reisizadeh et al. (2023) demonstrated that clipped SGD requires $\Omega(\epsilon^{-2})$ samples per iteration. Through this way, the constant arising from the variance of the approximation error can be controlled at the $\mathcal{O}(\epsilon^{2})$-level. This, in turn, relaxes the burden on iteration complexity, requiring only $\mathcal{O}(\epsilon^{-2})$ iterations for averaging during convergence proof. We leave technical discussions on optimizing the batch size while maintaining $\mathcal{O}(\epsilon^{-4})$ total sample complexity of IAN-SGD for future work.
Based on these assumptions, we first obtain the following upper bounds regarding $\|\nabla F(w_{t})\|$, $\|\nabla f_{\xi}(w_{t})\|$. See Proof 6 in Appendix G.
@@
Existing work (Zhang et al., 2019) studied the iteration complexity of clipped SGD (i.e., $\beta=1$ and $\xi^{\prime}=\xi$, without independence sampling) in the special bounded noise setting $\tau_{1}=0$, and obtained the complexity
$T\geq\Lambda\max\Big{\{}\frac{4\Lambda}{\epsilon^{4}},\frac{128L_{1}}{\epsilon},\frac{80L_{0}+512L_{1}\tau_{2}}{\epsilon^{2}}\Big{\}},$ (18)
@@
Another existing work Zhang et al. (2020a) analyzed a generalized clipping framework that combines momentum acceleration and gradient clipping, namely $w_{t+1}=w_{t}-\big{(}\nu\min\big{\{}c,\tfrac{\gamma}{\|m_{t}\|}\big{\}}m_{t}+(1-\nu)\min\big{\{}c,\tfrac{\gamma}{\|\nabla f_{\xi}(w_{t})\|}\big{\}}\nabla f_{\xi}(w_{t})\big{)},$ under a bounded noise setting (i.e., $\tau_{1}=0$). They establish a complexity of $\mathcal{O}\big{(}(F(w_{0})-F^{*})\tau_{2}^{2}L_{0}/\epsilon^{4}\big{)}$ when $\gamma/c=\Theta(\tau_{2})$, where $m_{t}$ denotes the momentum term and $c$ is the clipping threshold. This framework enables them to construct a recursion based directly on $\|\nabla F(w_{t})\|$, whereas our analysis is built around recursive bounds involving $\|\nabla F(w_{t})\|^{2}$ or $\|\nabla F(w_{t})\|^{2-\beta}$ at each step. Consequently, while both approaches yield the same order of convergence in terms of $\epsilon$, the complexity expressions slightly differ in the numerators due to different recursions.
Compared to the existing studies (Zhang et al., 2019; 2020b; Reisizadeh et al., 2023; Zhang et al., 2020a; Koloskova et al., 2023) on gradient clipping algorithms, Theorem 2 establishes convergence guarantees without the bounded approximation error assumption and the use of extremely large batch sizes that depend on the target error $\epsilon$. Through numerical experiments in Section 7 and ablation study in A.2 later, we show that it suffices to query a small number of independent samples for IAN-SGD in practice.